### v1

# Automated Contrastive Learning Strategy Search for Time Series

Abstract.
In recent years, Contrastive Learning (CL) has become a predominant representation learning paradigm for time series. Most existing methods in the literature focus on manually building specific Contrastive Learning Strategies (CLS) by human heuristics for certain datasets and tasks. However, manually developing CLS usually require excessive prior knowledge about the datasets and tasks, e.g., professional cognition of the medical time series in healthcare, as well as huge human labor and massive experiments to determine the detailed learning configurations. In this paper, we present an Automated Machine Learning (AutoML) practice at Microsoft, which automatically learns to contrastively learn representations for various time series datasets and tasks, namely Automated Contrastive Learning (AutoCL). We first construct a principled universal search space of size over $3\times 10^{12}$, covering data augmentation, embedding transformation, contrastive pair construction and contrastive losses. Further, we introduce an efficient reinforcement learning algorithm, which optimizes CLS from the performance on the validation tasks, to obtain more effective CLS within the space. Experimental results on various real-world tasks and datasets demonstrate that AutoCL could automatically find the suitable CLS for a given dataset and task. From the candidate CLS found by AutoCL on several public datasets/tasks, we compose a transferable Generally Good Strategy (GGS), which has a strong performance for other datasets. We also provide empirical analysis as a guidance for future design of CLS.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Time series is a sequence of data points recorded chronologically, which has been collected and analyzed for various real-world applications such as clinical diagnosis (Eldele et al., 2021; Zhang et al., 2022b), electricity usage forecasting (Zhou et al., 2021; Yue et al., 2022) and web traffic monitoring (Ren et al., 2019). A paramount challenge in the field of time series representation learning lies in developing an efficacious encoder capable of deriving informative embeddings. These embeddings should be readily applicable to a variety of downstream tasks, such as classification and forecasting, thereby enhancing the model’s utility and performance in predictive analytics.
A predominant representation learning paradigm is Contrastive Learning (CL) (Liu et al., 2021), of which the core idea is to train models by pulling the positive embedding pairs closer and pushing the negative embedding pairs far apart. CL has been widely explored in domains such as computer vision (Ozbulak et al., 2023), natural language processing (Zhang et al., 2022a) and graph mining (Wu et al., 2021). In recent years, CL has been introduced into the time series domain (Zhang et al., 2023). The preponderance of existing research in CL as applied to time series data primarily emphasizes the development of sophisticated data augmentation techniques and the construction of contrastive pairs. For instance, TS-TCC (Eldele et al., 2021) employs jittering and scaling techniques to produce varied perspectives of the input time series. These generated perspectives, originating from the same data sample, are then treated as positive instances within the model’s training regime. TS2Vec (Yue et al., 2022) uses random cropping to generate positive pairs and treat other data points in the context as negative samples. TF-C (Zhang et al., 2022b) augments the input via time-domain and frequency-domain masking, and embeddings from different samples are treated as negative pairs.
Despite the effectiveness of these approaches, the majority are tailored with manual construction of Contrastive Learning Strategies (CLS) specific to particular datasets and tasks, representing distinct examples within the broader spectrum of general CLS space, which could contain millions even trillions of choices. Crafting CLS for specific tasks or datasets in such a huge space through heuristic methods could be challenging, often demanding extensive domain expertise and considerable manual effort through numerous trial-and-error iterations. A natural question arises that is it possible to build a system to automatically find a suitable CLS for a given task? A pertinent question emerges: can we identify a CLS that is generally effective across a dynamic range of tasks and datasets?
Aiming at automating the process of machine learning, Automated Machine Learning (AutoML) (He et al., 2021) has become popular in recent years. In the field of CL, some works optimize data augmentation strategy on image (Tian et al., 2020) and graph (You et al., 2021) data. For time series, a recent work InfoTS (Luo et al., 2023) also searches the data augmentation based on information theory which yet does not consider the final goal of representation learning for downstream tasks. Moreover, these studies predominantly focus on data augmentation, they often overlook other crucial dimensions of CL, such as loss functions, the construction of contrastive pairs, and embedding transformations, which are typically crafted by human intervention. Specifically, a CLS encompasses diverse loss functions (e.g., InfoNCE (Oord et al., 2018) and Triplet Loss (Hoffer and Ailon, 2015)) and metric functions (e.g., dot product, cosine similarity, negative Euclidean distance) that affect representation learning quality (Liu et al., 2021). Contrastive pair construction also plays a key role to distinguish sample-level characteristics or temporal dependencies (Oord et al., 2018). Embedding techniques, including augmentation (Wang et al., 2019) and normalization (Wang and Isola, 2020), further influences the pretrained model performance. Few works have been conducted on jointly considering all these aspects of the contrastive learning on time series.
To reduce the burden of excessive prior domain knowledge and massive manual trials of CL on time series for specific tasks, in this paper, we present an Automated Contrastive Learning (AutoCL) framework at Microsoft, which aims at learning to contrastively learn the representations for various time series tasks. Specifically, we first construct a comprehensive solution space of CLS, which not only covers the most important configurable dimensions of CL, but also considers the modest range of options for each dimension. The whole space is comprised of dimensions including data augmentations, embedding transformations, contrastive pair construction and contrastive learning objectives, with up to $10^{12}$ level of options in total. To efficiently target the suitable CLS in such huge solution space, we further introduce a reinforcement learning algorithm which directly optimizes the CL performance on downstream tasks in validation datasets. Compared with the recent automation-based CL method InfoTS (Luo et al., 2023), our method takes the two-fold advantage from more comprehensive solution space and the direct optimization approach on CL, while InfoTS only searches data augmentation configurations based on task-agnostic criteria, which does not consider the intrinsic connection between CLS and the downstream task, thus may result in targeting an improper CLS.
We empirically evaluate our AutoCL framework and the derived CLS over three practical downstream tasks including classification, forecasting and anomaly detection on real-world time series datasets. We first directly apply AutoCL over each task to obtain its suitable CLS, and the results of which could significantly outperform existing CL works, demonstrating the superiority of AutoCL. Then we derive a Generally Good Strategy (GGS) based on the candidate CLS found by AutoCL from several tasks, and test GGS on all tasks. The results show that GGS could achieve remarkable performance for all tasks, showing its strong transferability across tasks and datasets. We also provide extensive empirical analysis on the relation between the candidate CLS and the model’s performance on downstream tasks, hoping to provide guidance for the future design of CLS. We also tested our derived solution in a real-world deployed application of epilepsy seizure detection task in Shanghai Ruijin hospital in China, which has demonstrated the general good performance on the realistic platform.
The major contributions of the paper are summarized as follows:
  • We introduce AutoCL to automatically derive suitable CLS for various tasks and datasets on time series. A comprehensive solution space covering the most critical dimensions of CL with an efficient reinforcement learning algorithm have been proposed.
  • Extensive experiments in both public benchmarks and a deployed application have demonstrated the superiority of the derived CLS with better performance compared to the existing CL methods.
  • From the candidate CLS, we first provide some empirical findings on CLS and downstream tasks to guide the future design CLS. We also obtain a Generally Good Strategy (GGS), which can be used as a strong baseline for various time series datasets and tasks.
2. Methodology
In this section, we first formulate the problem of the automated contrastive learning strategy search, and then introduce a method called AutoCL for the problem. We will elaborate the two components of AutoCL: the solution space and the search algorithm.
Figure 1. Illustration of Contrastive Learning Strategy (CLS).
2.1. Problem Formulation
Denote $\mathcal{D}$ as a time series dataset comprised of training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$ and test $\mathcal{D}_{\text{test}}$ sets; $\mathcal{T}$ as a downstream task along with evaluation metrics $\mathcal{M}$; and $f^{E}$ as the encoder which extracts embeddings of the input time series. The problem of automated Contrastive Learning Strategy (CLS) search has two sub-problem: (1) define a solution space $\mathcal{A}$; (2) build a search algorithm $f$ to find a suitable CLS $A\in\mathcal{A}$ based on the validation performance on $\mathcal{D}_{\text{val}}$ regarding $\mathcal{M}$, such that after pre-training with $A$ on $\mathcal{D}_{\text{train}}$, $f^{E}$ will have a good performance on $\mathcal{D}_{\text{test}}$.
2.2. Solution Space
Based on the studies in the literature and our own experiences, we use the following principles to define the space (You et al., 2020): (1) encompasses the key dimensions of CLS, incorporating insights from prior human-designed strategies; (2) accounts for a moderate spectrum of choices within each dimension.
An illustration of a CLS is shown in Figure 1. The essence of CL is to extract semantic-preserving embeddings invariant to small data perturbations (Tian et al., 2020), and thus data augmentation is the cornerstone dimension of CLS. Another commonly used practice to enhance the robustness of representation is embedding transformations, including embedding augmentation (Wang et al., 2019) and normalization. There are various aspects to construct positive and negative embeddings pairs, where different aspects reflect different characteristics of data. Different tasks might place emphasis on different aspects. Finally, different similarity metric functions and different types of contrastive learning losses could have divergent impact on the learned representations. In summary, we consider four dimensions: data augmentations, embedding transformations, contrastive pair construction and contrastive losses. A summary of the dimensions, sub-dimensions and their options is shown in Table 1, and we describe their details in the following content.
Table 1. The proposed solution space.
Dimensions Sub-Dimensions Options
Data Augmentations Resizing (Length) (Mehari and Strodthoff, 2022) $0.0,0.1,\cdots,0.9,0.95$
Rescaling (Amplitude) (Han et al., 2021)
Jittering (Eldele et al., 2021)
Point Masking (Zerveas et al., 2021)
Frequency Masking (Zhang et al., 2022b)
Random Cropping (Yue et al., 2022)
Order 0, 1, 2, 3, 4
Embedding Transformations Jittering (Wang et al., 2019) $0.0,0.1,\cdots,0.9,0.95$
Masking
Normalization None, LayerNorm (Ba et al., 2016), $l_{2}$
Contrastive Pair Construction Instance Contrast (Eldele et al., 2021) True
Temporal Contrast (Yue et al., 2022) True, False
Cross-Scale Contrast (Bachman et al., 2019) True, False
Kernel Size 0, 2, 3, 5
Pooling Operator Avg, Max
Use Adjacent Neighbor (Franceschi et al., 2019) True, False
Contrastive Losses Loss Types InfoNCE (Oord et al., 2018), Triplet (Hoffer and Ailon, 2015)
Similarity Functions Dot, Cosine, Distance
Temperature Parameters $10^{-2},10^{-1},10^{0},10^{1},10^{2}$
Data Augmentations. Data augmentations transform the input data into different but related views, which are the cornerstones of a CLS. We consider 6 most commonly adopted data augmentations as the sub-dimensions, including resizing (length) (Mehari and Strodthoff, 2022), rescaling (amplitude) (Han et al., 2021), jittering (Eldele et al., 2021), point masking (Zerveas et al., 2021), frequency masking (Zhang et al., 2022b), and random cropping (Yue et al., 2022). Each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be an input time series, where $T$ and $c$ are the temporal length and the number of variables. For point masking, $p$ refers to the ratio of the input data points to be masked: $\mathbf{x}=\mathbf{x}\odot\mathbf{m}$, where $\mathbf{m}[t]\sim Bernoulli(1-p)$. For jittering, $p$ is the standard deviation of the Gaussian noise: $\mathbf{x}=\mathbf{x}+\mathcal{N}(0,p)$. For more details of each augmentation, please refer to the corresponding paper. We discretize the value of $p$ into 11 choices: $[0.0,0.1,\cdots,0.8,0.9,0.95]$. For the first value, we use $0.0$ to represent the setting where a specific data augmentation is not enabled. For the last value, we do not use $1.0$ since it is meaningless for some data augmentations in practice. For example, in point masking, $p=1.0$ means masking out all the input data.
Additionally, the order of applying data augmentations also influences the learned embeddings (Cubuk et al., 2019). For instance, the outcome of applying point masking followed by random cropping differs from the sequence of first random cropping and then point masking. However, the ordering of other data augmentations, such as resizing and rescaling, does not have a significant influence on the data. Therefore, we design 5 different orders of applying data augmentations. For details, please refer to Appendix. In summary, there are totally $5\times 11^{6}$ options for data augmentations.
Embedding Transformations. Embedding augmentations (Wang et al., 2019) and normalization (Wang and Isola, 2020) have been explored and proven useful in the literature. For embedding augmentations, we consider embedding jittering (Wang et al., 2019) and embedding masking as the sub-dimensions, which are similar to the jittering and point masking in data augmentation operation. These two embedding augmentations are also parameterized by a parameter $p\in[0,1]$ whose value is also discretize $p$ into 11 values. For normalization, we consider to use no norms, $l_{2}$ norm, and layer norm (Ba et al., 2016) as the options. In summary, there are $3\times 11^{2}$ options for embedding transformations.
Contrastive Pair Construction. There are two types of embedding pairs: positive pairs, where embeddings are semantically related to each other, and negative pairs, where embeddings do not necessarily share hidden semantics. For time series, contrastive pairs can be constructed from three aspects: instance contrast (Eldele et al., 2021), temporal contrast (Yue et al., 2022), and cross-scale contrast (Bachman et al., 2019).
Let $\mathbf{h}_{1},\mathbf{h}_{2}\in\mathbb{R}^{T\times d}$ be the embeddings of the two views $\mathbf{x}_{1}$, $\mathbf{x}_{2}$ of an input $\mathbf{x}$, where $T$ and $d$ are the length and dimension size. The objective of the instance contrast is to distinguish embeddings of two different time series instances $\mathbf{x},\mathbf{x}^{\prime}$ by pulling $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$ (as well as $\mathbf{h}_{1}^{\prime}$ and $\mathbf{h}_{2}^{\prime}$) closer and pushing $\mathbf{h}_{1},\mathbf{h}_{2}$ far away from $\mathbf{h}_{1}^{\prime},\mathbf{h}_{2}^{\prime}$. The temporal contrast aims to preserve the local temporal information. For a time step $t$, it pulls $\mathbf{h}_{1}[t]$ and $\mathbf{h}_{2}[t]$ closer and optionally also $\mathbf{h}_{1}[t]$ and its temporally adjacent embeddings $\mathbf{h}_{1}[t-1],\mathbf{h}_{1}[t+1]$ closer. It pushes $\mathbf{h}_{1}[t]$ away from $\mathbf{h}_{1}[t^{\prime}]$, where $t^{\prime}\notin[t-1,t+1]$. The above two contrasts might be sufficient for short time series (e.g., $T<100$), however, they are incapable of capturing multi-scale dynamics of long time series (e.g., $T>1000$) (Yue et al., 2022). We follow (Yue et al., 2022) and apply hierarchical pooling over $\mathbf{h}_{1},\mathbf{h}_{2}$ to obtain multi-scale embeddings $\mathbf{h}_{1}^{(s)},\mathbf{h}_{2}^{(s)}\in\mathbb{R}^{T_{s}\times d}$, where $s\in[1,\cdots,S]$ is the scale index and $T_{s}$ is the length. The hierarchical pooling is associated with two parameters: pooling operator (average or max pooling) and kernel size ($[0,2,3,5]$). The above instance contrast and temporal contrast are applied for all the scales. In addition to the instance and temporal contrasts, cross-scale contrast (Bachman et al., 2019) could further improve the cross-scale consistency. Specifically, if $\mathbf{h}^{(s+1)}[t_{s+1}]$ is pooled from $\mathbf{h}^{(s)}[t_{s}]$, then these two embeddings should be pulled closer, otherwise, they should be pushed away from each other.
In practice, one must choose at least one contrastive aspect to perform CL. Based on the studies in the literature and our own experiences, the instance contrast is the most indispensable aspect, and therefore it is always included. As a result, there are 2=3-1 contrast aspects, 2 pooling operators, 4 kernel sizes and 2 options (whether to include temporal adjacent embeddings) for the temporal contrast, which means we have 32 options in total.
Contrastive Losses. Different types of contrastive losses behave differently in practice. For example, InfoNCE (Oord et al., 2018) guide the model’s learning by maximizing the probabilities of the positive pairs, and Triplet loss (Hoffer and Ailon, 2015) directly maximize the similarity score of the positive pairs. A contrastive loss, e.g., InfoNCE, is usually comprised of a similarity function, measuring the similarity between two embeddings, and a temperature parameter (Oord et al., 2018), controling the magnitude of similarly scores. For the similarity functions, we consider dot product, cosine and negative Euclidean distance. For the temperature parameter, we consider $[10^{-2},10^{-1},10^{0},10^{1},10^{2}]$. In summary, there are 2 types of losses, 3 similarity functions and 5 temperatures, resulting in 30 total options.
Figure 2. Illustration of the search algorithm of AutoCL.
2.3. Search Algorithm
As shown in Section 2.2, there are around $3\times 10^{12}$ total strategies in the solution space. Therefore, it could be extremely challenging and resource consuming to manually pick a suitable CLS from the entire space. In this paper, we aim to automate the process of searching the suitable CLS within the solution space. In the following content, we first present an overview of the algorithmic part of AutoCL, and then elaborate its details, including the controller network and the two phases: candidate search and candidate evaluation.
Overview. In recent years, Reinforcement Learning (RL) has become a popular paradigm to implement AutoML pipelines. Our proposed algorithm follows the RL paradigms in (Cubuk et al., 2019; Li et al., 2022; Zhao et al., 2021), and an illustration is shown in Figure 2. The search algorithm is comprised of two phases: candidate search and candidate evaluation. In the candidate search phase, the agent, i.e., controller network $f_{C}$, interacts with the environment, including encoder training and validation, etc., for a certain number of iterations. For each interaction iteration, $f_{C}$ first samples a CLS $A$ and send it to the environment. Then, the environment trains the encoder $f_{E}$ based on $A$ and returns a reward $R$ to the agent. Finally, the agent updates the controller network $f_{C}$ based on $R$. During the interactions, a CLS $A$ satisfying certain requirements is recorded in the candidate set $\mathcal{A}$. In the candidate evaluation phase, for each CLS candidate $A\in\mathcal{A}$, we train an encoder $f_{E}$ based on $A$, then evaluate $f_{E}$ on the downstream task, and record its performance on the validation data.
Controller Network. The controller network $f_{C}$ learns to sample CLS for the given dataset and downstream task. An illustration of the controller network is presented in Figure 3. The controller network is comprised of (1) a learnable embedding $\mathbf{e}\in\mathbb{R}^{d}$, which learns the information of the input dataset and task during training; (2) a project function, projecting $\mathbf{e}$ into a hidden space; (3) $N$ action branches, each of which samples an option (e.g., 0.5) for its corresponding sub-dimension (e.g., point masking) in the solution space. Formally, $a_{n}$ is sampled according to the probability $\mathbf{p}_{n}\in\mathbb{R}^{n_{K}}$:
(1) $\mathbf{p}_{n}=\text{MLP}_{n}(\sigma(\mathbf{W}\mathbf{e})),$
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ and $\sigma$ are the weight and the tanh activation of the project layer. We denote a full strategy as $A=[a_{1},a_{2},\dots,a_{N}]$.
Figure 3. Controller network $f_{C}$.
Candidate Search. During candidate search, parameters of both the controller $f_{C}$ and the encoder $f_{E}$ need to be updated. We denote the parameters of $f_{C}$ and $f_{E}$ as $\theta_{C}$ and $\theta_{E}$, respectively. Optimizing $\theta_{C}$ and $\theta_{E}$ is a bi-level optimization problem (Zhao et al., 2021; Li et al., 2022; Luo et al., 2023). Given the strategy sampled by the controller $A\sim f_{C}$, the lower-level problem is to find the optimal $\theta_{E}^{*}$ that minimizes the contrastive loss $L\in A$. The upper-level problem is to maximize the reward $R$ on the validation set by optimizing $\theta_{C}$. Formally, we have:
(2) $\centering\begin{split}\max_{\theta_{C}}\,&\mathbb{E}_{A\sim f_{C}}[R(A,\theta_{E}^{*})]\\ \text{s.t.}\,&\theta_{E}^{*}(A)=\arg\min_{\theta_{E}}L(\theta_{E},A).\end{split}\@add@centering$
In practice, optimizing $\theta_{C}$ after the completion of optimizing $\theta_{E}$, i.e., conducting the complete contrastive pre-training for $f_{E}$, could be extremely costly (Zhao et al., 2021; Li et al., 2022). Following (Liu et al., 2018), we use the first-order approximation of the gradient for the encoder $f_{E}$. Let $\eta$ be the learning rate in gradient decent, then $\theta_{E}^{*}(A)$ is approximated by:
(3) $\theta_{E}^{*}(A)\approx\theta_{E}-\eta\bigtriangledown_{\theta_{E}}L(\theta_{E},A)~{}.$
After obtaining $\theta_{E}^{*}(A)$ (or $f_{E}^{*}$), the next step is to obtain the reward $R(A,\theta_{E}^{*})$. We first combine $f_{E}^{*}$ with a downstream model $f_{D}$ to obtain the full model $f=f_{D}\circ f_{E}^{*}$. Next, we use the training data to train $f$ on the downstream task (e.g., classification), and we use the performance (e.g., accuracy) of $f$ on the validation data as the reward $R(A,\theta_{E}^{*})$. However, the raw reward $R$ cannot be directly used to update the parameters of the controller $\theta_{C}$, since for many metrics (e.g., accuracy or mean squared error), $R>0$ always holds true. Without a negative reward, the agent might be unable to learn meaningful behaviors (Williams, 1992). To address this problem, we use the maximum reward of previous $(n-1)$ steps $R^{*}=\max\{R_{1},\dots,R_{n-1}\}$ as the baseline for the $n$-th step. Therefore, the final reward $\Delta_{n}$ is:
(4) $\Delta_{n}=\alpha\cdot(R_{n}-R^{*}+\epsilon)$
where $\alpha>0$ is a scaling factor and $\epsilon>0$ is a small constant. Given $\Delta$, $\theta_{C}$ is updated based on the REINFORCE algorithm (Williams, 1992).
In the bi-level optimization, if the model in the lower-level optimization is trained along an undesired direction, it will take much more steps to bring it back to the desired direction (Li et al., 2022). To address this issue, in the lower-level optimization, rather than directly updating the parameters of the encoder $f_{E}$, we update the parameters of $\tilde{f}_{E}$, which is the copy of $f_{E}$. Let $\tilde{f}_{E}^{*}$ denote the updated encoder copy with parameters $\tilde{\theta}_{E}^{*}$. Then we can obtain the raw reward $R(A,\tilde{\theta}_{E}^{*})$ and the final reward $\Delta$ based on $\tilde{f}_{E}^{*}$. If $\Delta<0$, then we discard $\tilde{f}_{E}^{*}$ since it goes to the undesired direction. If $\Delta>0$, then we replace the original encoder $f_{E}$ with $\tilde{f}_{E}^{*}$, and also add $A$ to the candidate set $\mathcal{A}$, as the model is updated in the desired direction.
Candidate Evaluation. To evaluate each candidate $A\in\mathcal{A}$, we first pre-train the encoder $f_{E}$ on the training data based on $A$. Then we train a full model $f=f_{D}\circ f_{E}$ on the training data for the downstream task (e.g., classification), where $f_{D}$ is the downstream model. The evaluation score (e.g., accuracy) of $f$ on the validation data is recorded as the performance of $A$.
Complexity Analysis. Our proposed algorithm is efficient in both phases. In the phase 1, we only use the first order approximation rather than exhausted training; in the phase 2, we could take advantage of parallel computing. Denote $N_{cs}$ as the number of iterations in phase 1, $T_{ep}$ and $T_{val}$ as the time of training the encoder for one epoch, and the time of validation. The time complexity for the phase 1 is $O(N(T_{ep}+T_{val}))$. Denote $M$ as the number of iterations of the full contrastive pre-training, $|\mathcal{A}|$ as the total number of candidates, and $K$ as the number of machines available for evaluation, then the time complexity of the phase 2 is $O(\frac{|\mathcal{A}|}{K}(MT_{ep}+T_{val}))$. Therefore, the total complexity is $O((\frac{|\mathcal{A}|}{K}M+N_{cs})T_{ep}+(N_{cs}+1)T_{val})$ In comparison, if we do not use the first order approximation in Equation (3), then the complexity is $O(N_{cs}MT_{ep}+N_{cs}T_{val})))$ for phase 1, which is significantly larger than our current practice. Note that $N_{cs}\geq|\mathcal{A}|$.
Table 2. Classification results on HAR and Epilepsy.
Datasets Metrics AutoCL GGS InfoTS TS2Vec TS-TCC CPC Self-EEG
HAR ACC 0.963 0.937 0.930 0.930 0.904 0.839 0.653
F1 0.963 0.937 0.929 0.930 0.904 0.833 0.638
Epilepsy ACC 0.982 0.977 0.976 0.975 0.972 0.966 0.937
F1 0.972 0.963 0.962 0.962 0.955 0.944 0.892
Table 3. Anomaly detection results on Yahoo and KPI.
Yahoo KPI
Methods F1 Precision Recall F1 Precision Recall
SPOT 0.338 0.269 0.454 0.217 0.786 0.126
DSPOT 0.316 0.241 0.458 0.521 0.623 0.447
DONUT 0.026 0.013 0.825 0.347 0.371 0.326
SR 0.563 0.451 0.747 0.622 0.647 0.598
TS2Vec 0.745 0.729 0.762 0.677 0.929 0.533
InfoTS 0.746 0.744 0.747 0.672 0.927 0.526
GGS 0.755 0.766 0.745 0.689 0.807 0.601
AutoCL 0.758 0.808 0.715 0.694 0.901 0.565
Table 4. Univariate time series forecasting results on ETTh1/ETTh2/ETTm1. The lower, the better.
AutoCL GGS InfoTS TS2Vec CoST TNC TS-TCC
Dataset Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 24 0.036 0.142 0.038 0.146 0.039 0.149 0.039 0.151 0.040 0.152 0.057 0.184 0.103 0.237
48 0.052 0.173 0.054 0.177 0.056 0.179 0.062 0.189 0.060 0.186 0.094 0.239 0.139 0.279
168 0.087 0.223 0.088 0.227 0.100 0.239 0.142 0.291 0.097 0.236 0.171 0.329 0.253 0.408
336 0.103 0.247 0.103 0.248 0.117 0.264 0.160 0.316 0.112 0.258 0.192 0.357 0.155 0.318
720 0.113 0.266 0.123 0.274 0.141 0.302 0.179 0.345 0.148 0.306 0.235 0.408 0.190 0.337
ETTh2 24 0.076 0.204 0.076 0.205 0.081 0.215 0.091 0.230 0.079 0.207 0.097 0.238 0.239 0.391
48 0.106 0.246 0.107 0.248 0.115 0.261 0.124 0.274 0.118 0.259 0.131 0.281 0.260 0.405
168 0.169 0.314 0.172 0.318 0.171 0.327 0.198 0.355 0.189 0.339 0.197 0.354 0.291 0.420
336 0.184 0.344 0.188 0.346 0.183 0.341 0.205 0.364 0.206 0.360 0.207 0.366 0.336 0.453
720 0.189 0.354 0.196 0.359 0.194 0.357 0.208 0.371 0.214 0.371 0.207 0.370 0.362 0.472
ETTm1 24 0.013 0.084 0.014 0.085 0.014 0.087 0.016 0.093 0.015 0.088 0.019 0.103 0.089 0.228
48 0.024 0.114 0.025 0.117 0.025 0.117 0.028 0.126 0.025 0.117 0.036 0.142 0.134 0.280
96 0.036 0.142 0.038 0.147 0.036 0.142 0.045 0.162 0.038 0.147 0.054 0.178 0.159 0.305
288 0.071 0.201 0.078 0.211 0.071 0.200 0.095 0.235 0.077 0.209 0.098 0.244 0.204 0.327
672 0.100 0.240 0.113 0.255 0.102 0.240 0.142 0.290 0.113 0.257 0.136 0.290 0.206 0.354
Table 5. The Generally Good Strategy (GGS) derived by our AutoCL method.
Data Augmentations Embedding Transformations Contrastive Pair Construction Loss Functions
Resize Rescale Jitter Point Mask Freq. Mask Crop Order Emb. Jitter Emb. Mask Norm Instance Temporal Cross-Scale Kernel Pool Adj. Neighbor Loss Type Sim. Func. Scaling
0.2 0.3 0.0 0.2 0.0 0.2 3 0.7 0.1 None True False False 5 Avg False InfoNCE Distance 1.0
3. Experiments
In this section, we provide empirical evaluations to answer the questions: Q1. How well does the CLS found by AutoCL perform compared to existing methods? Q2. Does there exist a Generally Good Strategy (GGS) that has a strong performance on all settings? Q3. How does each component in AutoCL contribute to the final results? Q4. What can we learn from the candidates found by AutoCL? Q5. How does AutoCL perform in the real-world applications? We answer these questions in Section 3.2 - 3.6, respectively.
3.1. Experimental Setup
Datasets. We use 7 real-world public datasets and 1 private data. The public offline datasets can be grouped by three tasks: classification, anomaly detection and forecasting. For classification: we use HAR (Anguita et al., 2013), Epilepsy (Andrzejak et al., 2001; Eldele et al., 2021). HAR is comprised of 10,299 sensor readings collected from 30 subjects, where each reading corresponds to one of the 6 activities and has a length of 128. The number of samples in the training, validation and test data are 5,881/1,471/2,947. Epilepsy contains 11,500 electroencephalogram (EEG) time series segments in a length of 178 from 500 subjects, and the task is to recognize whether there is epileptic seizure. The train/validation/test split is 7,360/1,840/2,300. For anomaly detection, we use Yahoo (Laptev et al., 2015) and KPI (Ren et al., 2019). Yahoo has 367 hourly time series with human labeled anomaly points, which contains synthetic and real web traffic data to Yahoo! services. KPI is comprised of 58 minutely KPI curves collected from various Internet companies such as Tencent and eBay. The two anomaly detection datasets are originally split into 50%/50% for train/test. During searching, we further split the training data by 90%/10% for training and validation. For forecasting, we use ETTh1/h2/m1(Zhou et al., 2021), which are 2-year time series data of oil temperature of electricity transformers. ETTh1 and ETTh2 are hourly recordings, which contain 17,420 time steps in total, and ETTm1 is a 15 min recording, which contains 69,680 time steps. These datasets are split into 12/4/4 months for train/validation/test. The private online dataset contains SEEG (stereo-electroencephalogram) signals of 15 anonymized patients collected from a the epilepsy surgery department of a Chinese hospital - Ruijin Hospital. For each patient, 12-15 invasive electrodes each with 12-18 contacts is used to obtain about 3.5 hour SEEG readings. Two neurosurgery technicians jointly marked the onset and end times of the seizures. We follow (Chen et al., 2022; Andrzejak et al., 2001) to pre-process the dataset. The signal frequency is unified to 256Hz, and the time series are divided into segments length of 0.5s. The segments containing epileptic waves are considered as positive samples accounting for 1.6% of the data. Following (Andrzejak et al., 2001), we balance the two classes into 50%:50%, which results in 11,778 samples. We further split the data into 70%/10%/20% for train/validation/test for each patient.
Comparison Methods. (1) Automatic data augmentation learning method: InfoTS (Luo et al., 2023). (2) Handcrafted CL methods: InfoTS (Luo et al., 2023), TS2Vec (Yue et al., 2022), TS-TCC (Eldele et al., 2021), CoST (Woo et al., 2021), TNC (Tonekaboni et al., 2020), CPC (Oord et al., 2018) and Self-EEG (Sarkar and Etemad, 2020). (3) Anomaly detection methods: SR (Ren et al., 2019), DONUT (Xu et al., 2018), SPOT and DSPOT (Siffer et al., 2017). (4) We also include GGS (Section 3.3), which is a generally good strategy found by AutoCL.
Evaluations and Implementations. Please refer to Appendix C for evaluation metrics, baseline settings and implementation details.
3.2. Direct Application of AutoCL
In this subsection, we answer Q1 by demonstrating the effectiveness of AutoCL for searching the suitable CLS for a given dataset and task. The overall performance of different methods on 7 public benchmark datasets and 3 downstream tasks are presented in Table 2-4. Among all the baselines, InfoTS, which could automatically search the suitable data augmentations, performs better than other baselines, which use handcrafted CLS. This observation shows the effectiveness of automatically searching the data augmentations. AutoCL further outperforms InfoTS, indicating the overall effectiveness of the proposed search space and search algorithm.
3.3. Transferability Study
Although AutoCL could efficaciously discover the suitable CLS for given datasets/tasks, in many cases, searching could still be resource consuming. Therefore, we seek to find a Generally Good Strategy (GGS) (Q2) which can be used as a strong baseline and a good starting point for new datasets and tasks.
We search GGS on 3 datasets HAR/Yahoo/ETTh1, and then test it on all the datasets. The procedure of finding GGS is described as follows. (1) Collect sets of top $K$ candidate CLS from each of HAR/Yahoo/ETTh1, and use Cartesian product to obtain the $K^{3}$ pairs $(A_{1},A_{2},A_{3})$. (2) Obtain the top candidate pair by the number of the shared sub-dimensions across $A_{1},A_{2},A_{3}$, and add the shared options in GGS. (3) For the un-shared sub-dimensions, discard the options that lead to significant validation performance drop. For details, please refer to appendix.
Table 6. Ablation Study on HAR
Metrics AutoCL Data Aug. Only Full Pre-train $f_{E}$ w/o Reward Filtering
ACC 0.963 0.942 0.965 0.930
F1 0.963 0.941 0.964 0.931
Time (hours) 10.516 10.218 85.193 10.724
A GGS we found is shown in Table 5. There are some interesting findings for the 4 dimensions. For the data augmentations, small perturbations on the input time series are preferred. For the embedding transformations, large jitterring is preferred. Besides, as a compromise for generally good across different datasets and tasks, normalization is not adopted by GGS. In fact, we have observed in Section 3.5 that classification tasks could benefit from LayerNorm yet forecasting disfavors normalization. For the contrastive pair construction, the instance contrast is the most important while the other aspects, i.e., the temporal and cross-scale contrasts, are less important. For the contrastive losses, the popular InfoNCE loss is generally good for all datasets and tasks. For the similarity function, negative Euclidean distance is preferred over the popular dot product and cosine similarity. Euclidean distance measuress the distance between two point, yet cosine similarity and dot product also consider the angle between the two points. This observation indicates that distance is more effective than angles for embeddings of time series data.
The performance of GGS on all datasts/tasks is presented in Table 2-4. It can be observed that GGS outperforms all the baselines in general, but underperforms AutoCL. For practical usage, we leave the trade-off, between AutoCL from scratch (with better performance) and direct transferring GGS (with lower computational cost) to novel settings, to the readers.
(a) With reward filtering.
(b) Without reward filtering.
Figure 4. Training curves of the models.
3.4. Ablation Study
To answer Q3, we investigate the impact of different components of AutoCL on HAR, and the results are shown in Table 6.
Effectiveness of the solution space. Compared with existing studies, which only consider the data augmentation dimension of CLS, our proposed solution space is more comprehensive, which is comprised of 4 most important dimensions: data augmentations, embedding transformations, contrastive pair construction and contrastive losses (Table 1). We can observe that the CLS obtained from the full space significantly outperforms the CLS obtained from the data augmentation dimension alone.
Effectiveness of the first-order approximation and the separate search/evaluation phases. We compare AutoCL with the full pre-training of $f_{E}$ during phase 1, i.e., given the CLS $A$ sampled by $f_{C}$, we pre-train the encoder $f_{E}$ from scratch till the max number of iterations is reached. The results in Table 6 show that the CLS found by the first-order approximation (AutoCL) is competitive to the full pre-training of $f_{E}$ in terms of ACC and F1. However, AutoCL is more than 8$\times$ faster than the full pre-training. This is because AutoCL (1) only pre-trains $f_{E}$ for one epoch for each iteration in phase 1; (2) evaluates candidates in parallel in phase 2.
Effectiveness of reward filtering. In Equation (4), we use reward filtering to improve the process of RL. In Table 6, significant performance drop is observed if we remove the reward filtering. In Figure 3(a), we can see that AutoCL can progressively obtain higher and higher validation ACC. However, as shown in Figure 3(b), if we remove reward filtering, the model fails quickly after the initial exploration in the first 100 steps.
3.5. Empirical Analysis of the Candidates
To answer Q4, we provide in-depth empirical analysis based on the candidate CLS, and hope to provide potential insights and guidelines for future CLS designs. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1 are shown in Figure 5, where rows correspond to datasets and columns correspond to sub-dimensions.
Loss Types. The 1st column shows that InfoNCE is generally better than Triplet, indicating that comparing the similarities scores of the positive/negative pairs in the probability space usually yields better performance than comparing them in their original space.
Similarity Functions. The 2nd column shows that dot product and negative Euclidean distance are generally better than cosine similarity. We believe this is because cosine similarity only captures the angle between two data points, yet ignores the magnitudes.
Temporal Contrast. The 3rd column indicates that the temporal contrast is generally good for forecasting and anomaly detection, but it is less useful for classification. During pre-training, the input time series of ETTh1 and Yahoo are very long (2,000 steps) but the inputs of HAR are very short (128 steps). This observation shows that the temporal contrast is more valuable for long time series.
Normalization. The 4th column shows that LayerNorm is important for classification, however, the embedding normalization has negative impact for forecasting and anomaly detection. We suspect this is because the normalization might destroy the fine-grained information, e.g., small anomaly points, which might be critical for forecasting and anomaly detection.
Embedding Jittering. The 5th column indicates that applying jittering on embeddings is generally useful. This finding is consistent with (Wang et al., 2019) that jittering injects various small semantic perturbations into the embeddings, which helps the model to capture the core semantics during contrastive pre-training.
Frequency Masking. In the last column, Figure 4(r) shows that frequency masking with high masking ratio could easily lead to good performance (low MSE). We believe this is because frequency masking could remove the short-range noisy patterns, and make the model focus on the long-range patterns. However, Figure 4(l) shows that the frequency masking has negative impact on anomaly detection. Essentially, the anomaly points are sudden and short-range noise, which only accounts for a small ratio, e.g., 1% of the entire time series. Fourier Transform (FT) and Inverse Fourier Transform (IFT) involved in the frequency masking can easily remove these points from the input, and thus the model is unable to learn the information about anomaly points during pre-training.
Table 7. Classification results in deployed application.
Metrics AutoCL GGS+LayerNorm GGS InfoTS TS2Vec
ACC 0.760 0.739 0.737 0.735 0.728
F1 0.756 0.734 0.730 0.730 0.719
3.6. Experiments in Deployed Application
To answer Q5, we deploy AutoCL to the machine learning platform of our partner hospital. The medical practitioners can easily utilize our AutoCL to obtain the effective time series models for analyzing various physiological time series data collected in medicine, e.g., EEG and SEEG time series signals in epilepsy healthcare.
We compare AutoCL and GGS with strong baselines, i.e., InfoTS and TS2Vec, on detecting epileptic seizure from SEEG data. These experiments, conducted with anonymized data, have received ethical approval from the hospital. The results are shown in Table 7. First, AutoCL has the best overall performance, showing the superiority of AutoCL in real-world applications. Second, GGS is slightly better than the recent automatic data augmentation method InfoTS, indicating that GGS has a strong transferability and is a strong baseline for new datasets and tasks. Consequently, in scenarios where a comprehensive search is prohibitively extensive, GGS presents a viable alternative. Moreover, empirical evidence from Section 3.5 suggests that the incorporation of Layer Normalization typically enhances performance. Integrating Layer Normalization within the GGS, as evidenced by that in Table 7, yields performance improvements, thereby affirming the efficacy of this insight.
(a) HAR: Loss Types
(b) HAR: Sim. Functs.
(c) HAR: Temporal
(d) HAR: Norm
(e) HAR: Emb. Jitter
(f) HAR: Freq. Mask
(g) Yahoo: Loss Types
(h) Yahoo: Sim. Functs.
(i) Yahoo: Temporal
(j) Yahoo: Norm
(k) Yahoo: Emb. Jitter
(l) Yahoo: Freq. Mask
(m) ETTh1: Loss Types
(n) ETTh1: Sim. Functs.
(o) ETTh1: Temporal
(p) ETTh1: Norm
(q) ETTh1: Emb. Jitter
(r) ETTh1: Freq. Mask
Figure 5. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1. X-axis: options. Y-axis: ACC/F1/MSE(48 horizon). ACC/F1: the higher the better. MSE: the lower the better. Low, Medium and High correspond to $0.1\sim 0.3$, $0.4\sim 0.6$ and $0.7\sim 0.95$ for embedding jittering and frequency masking ratio.
4. Related Work
Time Series Contrastive Learning. Contrastive learning aims to train encoders that can learn informative embeddings by distinguishing positive data pairs and negative data pairs (Liu et al., 2021), and it has been demonstrated to be effective in domains such as computer vision (Ozbulak et al., 2023), natural language processing (Zhang et al., 2022a) and graph mining (Liu et al., 2022). In recent years, some studies introduce contrasive learning into the time series domain and have achieved impressive performance (Zhang et al., 2023). SRL (Franceschi et al., 2019) proposes to treat two adjacent sub-sequences from the same time series as positive pairs, and sub-sequences from other time series as negative samples. CPC (Oord et al., 2018) introduces the popular InfoNCE loss. Self-EEG (Sarkar and Etemad, 2020) augments the input EEG signals with various transformations. TNC (Tonekaboni et al., 2020) constructs contrastive pairs based on the local smoothness of time series. TS-TCC (Eldele et al., 2021) transforms the input data with weak and strong augmentations, and use temporal and contextual contrasting to train the encoders. Instead of only considering data augmentations in the temporal domain, CoST (Woo et al., 2021), BTSF (Yang and Hong, 2022) and TF-C (Zhang et al., 2022b) further propose to augment time series in the frequency domain. TS2Vec (Yue et al., 2022) obtains time series representations of multiple scales by hierarchical pooling and performs CL at each scale. To reduce human efforts in finding the optimal data augmentations, InfoTS (Luo et al., 2023) recently proposes to automatically select data augmentations through meta-learning. However, InfoTS only considers data augmentations yet ignores other critical components in contrastive learning such as embedding transformation, contrasive pair construction and loss functions. Besides, InfoTS selects data augmentations based on a dataset/task-agnostic criteria, which is less effective in capturing the characteristics of the datasets and downstream tasks and thus might not be able find the optimal solution. Compared with InfoTS, AutoCL  (1) considers a larger search space comprised of data augmentations, emebedding transformations, and loss functions; (2) finds the optimal strategy based on the model’s performance on the validation set.
Automated Contrastive Learning. Despite various powerful contrastive strategies have been developed for particular datasets and downstream tasks, they cannot be easily adapted to other datasets and tasks (Luo et al., 2023; Tian et al., 2020; You et al., 2021). Therefore, it is necessary to develop automatic searching algorithms to find the optimal contrastive strategies for a given dataset and task. In the computer vision domain, InfoMin (Tian et al., 2020) introduces a mutual information minimization criteria for selecting augmented views. SelfAugment (Reed et al., 2021) uses the self-supervised image rotation task to guide the selection of data augmentations. In the graph domain, JOAO (You et al., 2021) presents a Bayesian framework to adaptively find the optimal data augmentations. AutoSSL (Jin et al., 2021) learns the weights of different contrastive strategies. AutoGCL (Yin et al., 2022) builds an end-to-end framework to automatically learn the distributions of node level augmentations. In the time series domain, InfoTS (Luo et al., 2023) recently propose to adaptively select the optimal data augmentations based on information theory. Different from previous studies which mainly focus on searching the optimal data augmentations, in this paper, we propose AutoCL to simultaneously consider various dimensions of CL, including data augmentations, embedding transformations, contrastive pair construction and loss functions.
5. Conclusion and Future Work
In this study, we introduced an innovative framework, AutoCL, for the automated discovery of Contrastive Learning Strategies (CLS) in time series data, aimed at diminishing the reliance on extensive domain expertise and iterative experimentation in the development of CLS tailored to specific datasets and tasks. The AutoCL framework is structured around two principal components: a comprehensive solution space, meticulously designed to encapsulate four critical dimensions of CLS, and an effective optimization algorithm to identify suitable and transferable CLS configurations. Our empirical tests on eight datasets demonstrate AutoCL’s effectiveness in finding suitable CLS, with the derived Generally Good Strategy (GGS) showing robust performance across different tasks. This study not only showcases the potential of AutoCL but also provides valuable insights for future contrastive learning research.
Given the versatility of our automation approach, we are confident that our proposed framework can be adapted to non-contrastive self-supervised learning approaches involving time series data, an avenue we intend to explore in our future research.
Appendix A Data Augmentations
Data augmentations transforms the input data into different but related views, which are the cornerstones of a CLS. In general, each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be the input time series, where $T$ and $c$ are the length and the number of variables, then the details of data augmentations are given below.
Resizing (length)
$p$ refers to the standard deviation of the Gaussian noise. The final length is given by
Rescaling (amplitude)
$p$ refers to the standard deviation of the Gaussian noise. The final amplitude is given by
Jittering
$p$ refers to the standard deviation of the Gaussian noise: $\mathbf{x}=\mathbf{x}+\mathcal{N}(0,p)$.
Point Masking
$p$ refers to the ratio of the input data points to be masked: $\mathbf{x}=\mathbf{x}\odot\mathbf{m}$, where $\mathbf{m}[t]\sim Bernoulli(1-p)$.
Frequency Masking
$p$ refers to the ratio of the frequencies to be masked. We first apply Discrete Fourier Transform (DFT) over $\mathbf{x}$ to obtain its frequency representation $\mathbf{x}_{f}=DFT(\mathbf{x})$. Next, we randomly mask out $p$ frequencies $\mathbf{x}_{f}=\mathbf{x}_{f}\odot\mathbf{m}$, where $\mathbf{m}[t]\sim Bernoulli(1-p)$. Finally, we transform $\mathbf{x}_{f}$ back to the time domain via Inverse Discrete Fourier Transform (IDFT): $\mathbf{x}=IDFT(\mathbf{x}_{f})$.
Random Cropping
$p=\frac{T^{\prime}}{T}$, where $T$ and $T^{\prime}$ are the lengths of the input $\mathbf{x}$, and the length of the common segment shared by two cropped sub-sequences $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. Specifically, $\mathbf{x}$ is cropped into two sub-sequences with a shared segment, i.e., $\mathbf{x}_{1}=\mathbf{x}[t_{1}:t_{1}^{\prime}]$ and $\mathbf{x}_{2}=\mathbf{x}[t_{2}:t_{2}^{\prime}]$, where $t_{1}\leq t_{2}<t_{1}^{\prime}\leq t_{2}^{\prime}$ and $\mathbf{x}_{c}=\mathbf{x}[t_{2}:t_{1}^{\prime}]$ is shared by ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$. Then $T^{\prime}=t_{1}^{\prime}-t_{2}$ and $p=\frac{T^{\prime}}{T}$.
Orders of Data Augmentations
The order of applying data augmentations also influences the learned embeddings (Cubuk et al., 2019). For instance, the outcome of applying point masking followed by random cropping differs from the sequence of first random cropping and then point masking. However, the ordering of other data augmentations, such as resizing and rescaling, does not have a significant influence on the data. Therefore, we design 5 different orders of applying data augmentations. The details are presented in Table 8.
ID. Order-1 Order-2 Order-3 Order-4 Order-5
1 Resize Resize Resize Resize Resize
2 Rescale Rescale Rescale Rescale Crop
3 Freq. Mask Freq. Mask Freq. Mask Crop Rescale
4 Jitter Jitter Crop Freq. Mask Freq. Mask
5 Point Mask Crop Jitter Jitter Jitter
6 Crop Point Mask Point Mask Point Mask Point Mask
Table 8. Orders of data augmentations.
Appendix B Procedure of Composing GGS
In this section, we elaborate the details for the procedure of composing GGS.
First, collect top 10 candidate CLS sets $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, $\mathcal{A}_{3}$ from HAR, ETTh1 and Yahoo, and construct candidate pairs: $(A_{1,i},A_{2,j},A_{3,k})$, where $A_{1,i}\in\mathcal{A}_{1}$, $A_{2,j}\in\mathcal{A}_{2}$, $A_{3,k}\in\mathcal{A}_{3}$.
Second, rank the candidate pairs by the number of their shared options. Note that for a probability option (e.g., jittering) $p\in[0.0,0.1,\cdots,0.9,0.95]$, it will be considered as a shared option, if $(\max(p_{1},p_{2},p_{3})-\min(p_{1},p_{2},p_{3}))<0.35$, where $p_{1}$, $p_{2}$, $p_{3}$ correspond to $A_{1,i},A_{2,j},A_{3,k}$.
Third, pick the pair that has the most shared options, and add their shared options in GGS. For the rest un-shared options, we search their values based on heuristics. The strategies with significant validation performance drop on any of HAR/ETTh1/Yahoo are discarded. We stop searching when a strategy with acceptable validation performance is found.
Appendix C More Experimental Setups
C.1. Evaluation Metrics
Our evaluation follows (Yue et al., 2022). For each dataset, we first pre-train an encoder based on a CLS, and then train a downstream model, which is evaluated on the test data. For classification, the downstream model is the SVM classifier with RBF kernel, and the evaluation metrics are accuracy (ACC) and F1. For forecasting, the linear regression model with $l_{2}$ norm is used, and the evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE). For anomaly detection, we follow the protocol used by (Yue et al., 2022; Ren et al., 2019). The evaluation metrics are F1, precision and recall.
C.2. Baselines
We consider the following baselines: InfoTS (Luo et al., 2023) adaptively learns optimal augmentations. TS2Vec (Yue et al., 2022) performs contrastive learning in an hierarchical manner. TS-TCC (Eldele et al., 2021) uses temporal and contextual contrasting to train encoders. CoST (Woo et al., 2021) performs augmentations in both temporal and frequency domains. TNC (Tonekaboni et al., 2020) treats adjacent samples as positive pairs and non-adjacent samples as negative pairs. CPC (Oord et al., 2018) introduces InfoNCE. Self-EEG (Sarkar and Etemad, 2020) uses various transformations to augment the input time series. SR (Ren et al., 2019) detects anomaly points based on spectral residual. DONUT (Xu et al., 2018) detect anomalies based on Variational Auto-Encoder (VAE) (Doersch, 2016). SPOT and DSPOT (Siffer et al., 2017) use the extreme value theory (Beirlant et al., 2006) to find anomalies. Note that GGS (Section 3.3) is the generally good strategy found by AutoCL, which has strong performance on different datasets and tasks.
C.3. Implementation Details
For the controller network, its hidden dimension size is 320, and MLPs are linear layers with a softmax activation. We use a 10 layer dilated CNN as the encoder similar to (Yue et al., 2022), and the sizes of hidden dimension and the output embeddings are 64 and 320. The learning rate of the controller is 0.0001. We fix the maximum number of the iterations in phase 1 as 500, and $\alpha$ as 10. For classification, forecasting and anomaly detection, the tolerance $\epsilon$ are 0.001, 0.0001 and 0.001, and the validation metrics are ACC, MSE (48 horizon) and F1. In pre-training, we fix the maximum length of the input time series as 2,000.